MULTISPECTRUM ANALYSIS OF THE ν_{4} AND ν_{2} BANDS OF ^{12}CH_{4}

\underline{MARY ANN H. SMITH}, Science Directorate, NASA Langley Research Center, MS 401A, Hampton, VA 23681-2199; D. CHRIS BENNER and V. MALATHY DEVI, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795.

Self- and air-broadened halfwidth and pressure-induced shift coefficients and their temperature dependences have been determined for a number of transitions in the ν_{4} and ν_{2} bands of ^{12}CH_{4} from laboratory absorption spectra recorded at room temperature and below with the McMath-Pierce Fourier transform spectrometer of the National Solar Observatory. In addition, accurate line center positions and absolute intensities were determined. The results were obtained by using a multispectrum nonlinear least squares technique^{a} to fit simultaneously 20 or more spectra recorded at high resolution (0.006-0.01 cm^{−1}) with path lengths between ∼ 1 and 150 cm and sample temperatures ranging from 210 to 314 K. For both self- and air broadening in the J-manifolds of the P and R branches of the ν_{4} band, we observed line mixing between certain transitions within the same manifold. In these cases the off-diagonal relaxation matrix elements were determined in the fits. Most of the air-broadened widths of unmixed ν_{4} lines retrieved using the multispectrum fits agree well with earlier values determined from single-spectrum fits,^{b} and the multispectrum results have smaller statistical uncertainties. The inclusion of line mixing in the fits was seen to have a greater effect on the retrieved values of the line shifts than on the retrieved values of other parameters.

aD. Chris Benner et al., JQSRT \underline{53}, 705-721 (1995).

bM. A. H. Smith et al., Spectrochimica Acta \underline{48A}, 1257-1272 (1992).